OpenStudy (anonymous):

find d/dx of 7x^2(x-6)

OpenStudy (anonymous):

i know answer is 7x^2(x-9) all over 3

OpenStudy (anonymous):

i know 7x^2 dx= 7x^3/3 so not sure why it is different now

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ d }{ dx } 7x ^{2} = 14x\] \[\frac{ d }{ dx } ( x - 6 ) = 1\] \[\frac{ d }{ dx } 7x ^{2}( x - 6 ) = 7x ^{2} + ( x - 6 ) (14x) = 21x ^{2} - 84x \]

OpenStudy (anonymous):

sorry i want the integral